Determine a, h and k, of the translated function g(x) based from the description, then write the equation of the quadratic function in vertex form.
1.) Vertex: (2,0); point (1,4)
solution:
2.) Vertex: (-3,6); point(-5,2)
solution:

Answer by mananth(16946)

The equation in vertex form is
y=a(x-h)^2+k
Vertex :(2,0); point (1,4)
y= a(x-2)^2+0
plug values of the point +0
4 = a(1-2)^2+0
4= a(1)
a=4
y = 4(x-2)^2
.